3. Suppose that the amount of liquid in a bottle is normally distributed with a mean of 16 oz and a standard deviation of 0.2 oz. If we take a random sample of 10 bottles from a production run, what is the probability that the sample mean is between 15.99 and 16.01?
Answer #1

Solution :

Given that,

mean = = 16

standard deviation = = 0.2

= / n = 0.2 / 10 = 0.0632

= P[(15.99 - 16) / 0.0632< ( - ) / < (16.01 - 16) / 0.0632)]

= P(-0.16 < Z < 0.16)

= P(Z < 0.16) - P(Z < -0.16)

= 0.5636 - 0.4364

= 0.1272

Probability = 0.1272